The length of a rectangle garden is 6 feet longer than its width.if the area is 55 sq.ft.find the length and width. Quadratic equation
Answer by checkley77(12844)

L=W+6
LW=55
(W+6)W=55
W^2+6W-55=0
(W+11)(W-5)=0
W-5=0
W=5 ANS. FOR THE WIDTH.
L=5+6=11 ANS. FOR THE LENGTH.
PROOF:
5*11=55
55=55